please help.
solve the system by substitution.
5x+4y=-2
y=4x+10

20x-4y=-8
y=2+5x

Answer by rapaljer(4671)

5x+4y=-2
y=4x+10

5x+4y = -2
5x+4(4x+10) = -2
5x+16x + 40=-2
21x =-2-40
21x=-42
x=-2
y=4x+10
y=4(-2)+10
y=-8+10
y=2

Check in the other equation:
5x+4y=-2
5(-2)+4(2)=-2
-10+8=-2 It checks!!

Now, you solve your other problem just like this one!!

R^2 at SCC